Scraps of text for incorporation in the VIS Meeting Minutes
- SR3: With new geophone inter-calibration factors, Lucia implemented the inertial damping and updated Guardian.
- Lucia implemented inertial damping and upated Guardian.
- Fabián wrote a Jupyter notebook based on Terrence's tutorial for blending geophone and LVDT using the methods he developed. It's still a work on progress.
- PRM: Ushiba-san did diagonalization of the BF and oplev.
- IP LVDT and geophone diagonalization, actuation diagonalization.
- Inter-calibration between LVDTs, geophones and ground seismometer.
- Sensor correction with an improvement of 2.5 times in the integrated rms.
- A bump appeared in the IP geophone TFs, very much like in SR2 in the past. New geophone calibration TFs were calculated, implemented in filter banks and are working fine.
- Inter-calibration between geophones and LVDTs were calculated again but not yet written in filter banks.
- It seems the resonance in IP-L appears at different frequencies, depending whether it's measured with the force TF or the IP free-swing ASD (66 and 80 mHz respectively). Double-checking this seems necessary.
- I removed the sensor and actuation diagonalization matrices that I had implemented last week because they introduced large coupling at some frequencies. Nevertheless, Lucia realized that diagonalization of only actuation is beneficial.
- I implemented inter-calibration with the seismometer on the ground and then implemented the LVDT sensor correction. With the microseimic conditions at the time, the seismoc noise in the LVDT decreased by 5.8 and 6.6 in L and T respectively, without increasing the rms.
- Beam splitter:
- Lucia tested the blending of the geophones with the corrected LVDTs,and the inertial dampiong. It seems to be working fine.
- Ushiba-san implemented low noise control for the payload.
- Lucia implemented and tested strong damping control in PRM.
- In BS, the IM picomotor cable was checked becuase pin 8 has a connection to ground. It was determined that the problem is deep within the chamber and we cannot fix it from outside. The yaw picomotor should still be functional because such a pin is connected to ground at the driver.
- In SR suspensions, we withdrew security stop screws to allow larger motion of the wholew suspenion in the longitudinal direction by approximately 5 mm. Clearly, we didn't move the stop screws around the F1 filters.
- We installed dial gauges in all four Type-B chambers underneath the mushroom heads. The aim is to monitor the motion of the chamber when pump down happens.
- We took pictures of all the security structure stop screws.
- Beam splitter cable health check was done again.
- The alignment of the SR3 oplev was modified to jave the beam far away from the lower breadboard. The aligmment of the Y arm was checked.
- Miyoki-san recovered the O3GK temperature of BS, and the fishing rod and mass on the F0 keystone was adjusted.
- The height of the four hanging optics were checked to be close enough to the desired values.
- I compared the BS IM transfer function with the SUMCOM model. The motivation was that, apparently the BS measurement had more peaks than the SR ones. Nevertheless, nothing wrong was found, the SUMCOM model and the measurements seem to be consistent. I used a list of resonant modes, but I shoud generate the calculated plots and show them along the measured ones.
- The BF Damper actuation was offloaded to the Traversers in all suspensions.
- We checked the polarity of the magnets in the BF Dampers and IMs in all suspensions. The TM magnets are not accesible.
- We took pictures of all the security structure screws.
- We recalibrated SR2 SF LVDT.
- Yamamoto-kun write a script to calculate the IP diagonalization matrices from the geometrical information. In some cases it coincides with what we had, althoug in others it was different because people carried out additional diagonalization. I changed all the matrices to the geometrical ones in order to check the TF in such conditions. The new TFs were posted in the klog this morning.
- Hirata-san corrected some discrepancies in some dimensions in the IP 2D-CAD.
- We checked SR3 V2 and SRM H1 OSEMs becuase the individual TFs had different amplitudes than in other suspensions. In both cases, we exchanged cables with other OSEMs and measured TFs again. In both cases the origin of the discrepancy seems to come from within the vacuum chamber.
- Ushiba-san plotted most of Type-B transfer functions in a way suitable for comparison. After remeasurig a few of them, I have the impression that the suspensions are healthy, although it seems there's still an unknown: in BS IM-R and IM-P see to have additional peaks in in SRs. I need to check in more detail.
- We measured TFs at the individual coils and plotted them in a suitable way for comparison. I should double-check, but a cursory comparison suggessts they are all fine. No klog yet.
- We checked the polarities of all the coil-magnet actuators in the four suspensions and everything seems consistent with what we know. No klog yet.
- On request from Takahashi-san: the oldest PR2 TF measurements I was able to find were from 26th of July 2019.
- In SR3, we removed the First Contact on Monday and released the suspension. Then AOS team installed the midsize baffle in SR3 and SRM.
- Yokozawa-san and Fabian measured the beam profile and power of the SR2 oplev.
- In PRM, we calibrated again the BF Damper vertical LVDTs by moving the BF with the SF keystone. The new values were written in the medm screen and accepted in the SDF.
- In PRM, We calibrated again the BF Damper horizontal LVDTs using the oplev and the design value of the position of the LVDTs. There is an apparent inconsistency with the sign and I need to double-check. I haven't written the new values in the medm screen yet.
- We did cable health check in the three PR suspensions.
We inspected SRM payload and chamber and we think we can exchange the small mirror after the installation of the mid-size baffle. When are we going to change it?
- In SR3, we locked the suspension and First Contact was applied.
- In the corresponding filter bank, I updated the information of the transfer function of the new geophone.
- In PR3 and PR2, we tested the Traverser limit switches.
- In PRM, OSEMs H2 and H3 were close to the end of their linear ranges, therefore, we moved them. The payload is not as stable as we would've liked, and was disturbed when we installed the Traverser limit switches. See klog.
- In PR3, we removed the First Contact on Monday ad released the suspension. As expected, by locking it for cleaning, we disturbed the pitch of the mirror. However the misaligment this time was different than others before (see a klog from today).
- Ikeda-san and Hirata-san did in-air cabling for the Traverser.
- In PR2 we installed the Traverser limit switches and new cable clamps. A few cables broke in the process but they were fixed. We also did in-vacuum and in-air cabling for the Traverser.
- Beam Splitter: after First Contact work last week, the H2 OSEM readout was close to the end of its usable range.
- We found a cable from one of the RM coils touching the security structure. We moved it and clamped it in a slightly different position.
- Then we adjusted the pitch and yaw of the BS using the picomotor and F0 yaw stepper motor respectively.
- We also inspected that nothing around the BF was touching the security structure. Ushiba-san also made a quick test moving up and down the F0 and checking the output of F1 and BF LVDTs.
- Finally, we moved H2 and H3 OSEMs to more suitable positions.
- We measured TFs and corroborated that the system is healthy.
- In SR2, the geophones stopped working, so we checked. The problem was tracked down to the geophone #2 pins powering the preamplfier. There is a short circuit between pins 4 and 5. Next week we will check the in-vacuum cables and, depending on the result, we may want to change the preamplfier or the whole geophone with its pod.
- Takahashi-san compared SR2 IP TF measurements with those of SRM, and found some discrepancies. The problem was likely produced by the geophone integrator being off. In SR2 it has to be turned on by hand. I measured the same TFs this week and they are fine.
- With the aim of understanding Terrece's work, in SR2 I tried to do the inter calibration between geophoes and LVDT. At low frequencies the inter calibration seems be good already, but around 1 Hz, where the resonant frequency of the geophone is, there is a deformation of the TF. I've been looking into this, but I don't have anything relevant to report yet.
- We replaced a broken LVDT in PRM. Namey, the one for the BF Damper horizontal LVDTs:
- In the old board we measured the resitance values determining two gains, and copied them to the new board.
- After installing the new board, we adjusted the phases in order to achive maximum ouput.
- Then, in order to calculate the new LVDT calibration factors, we measured the transfer functions from Y actuation to H1, H2 and H3 outputs, and compared with previous measurements.
- The DC amplitudes and phases of transfer functions are the same as before.
- Hirata-san continued working in the design of the traverser limit switches. Differences between the real hardware and the 3D-CAD have made the process difficult, but he's now working on a final version. He has reported his work at the ATC-KAGRA meeting.
- Takahashi-san found out that the LVDT drivers used in PR3 and PRM are using a reference signal with an amplitude of 12 Vpp.
- Together with Takahashi-san, we measured the whole set of relevant transfer functions in SR2.
- We compared with measurements from July 2019 and the differences were not relevant. The behaviour of the suspension seems very stable.
- We recorded in the klog the DC amplitudes of all transfer functions for the purposes of health check.
- Today Friday I might have disturbed the the oplev.
- Later, It was possible to engage ALIGNED state with Guardian properly, so it seems fine. Still, we should measure the mirror stage transfer functions on Monday.
- The readout of the QPDs was low (a few 100s of counts per segment), so I changed he gain in the QPD boards to high with the on-board switch.
- In the BS payload, the moved the security structure screws closer to their corresponding bodies. The gaps are around 3 or 4 mm.
- The BS RM coil cables had some plastic sheets that we should have removed after installation but we forgot. We removed them today.
- Hirata-san is still considering which one of the two methods of holding the traverser limit switches is more suitable. In the method which seems more convenient, the switches assembly would be pushed with 28 kgf (one motor considered) in case the switch malfunctions.
- We recovered PR3 pitch alignment after the earthquake. We didn't have to move OSEM V1.
- In PR3 oplev, I measured the distances between the optical components.
Miyo-kun sorted out bugs affecting Type-B automeasurement script. There are still some suggestions to tell him.
- We found out the channel name assigment of the BF LVDTs. There was no information about it. We disconnected in-vacuum cables and checked which channel in the medm screen was affected. Hirata-san wrote the information in a document.
- We inspected PR2 and PRM traversers to learn their orientations with respect to the suspensions. Hirata-san and others compiled the information in a document.
- We tested the two ways of holding the traverser limit swtiches. Both methods seem good, although one has the disadvantage that some set screws must be used at the further side of the assembly, and they may be accidentally dropped on the suspension.
For the BS:
- We inspected visually the BS payload and the BF in order to make sure no part of the suspension was too close to the security structure. It was all clear. Actually, the most of the security structure stop screws were too far away from the suspension. We brought closer a few of them hut others remain too far. We should move them to a more suitable place in the near future.
- The cable of the F0 fishing rod had a problem that we fixed. Probably, the origin of the problem was a loose screw in a connector.
- I gave information to Miyo-kun for the transfer function auto-measurement system.
In PRM, we tested the prototype of the cable clamps for the uppermost stage. The assembly works well but modifications are necessary:
- More places for cables are needed. We need at least nine and the prototype has only five.
- We need to devise a way to keep the cable shield, that is exposed, from touching the clamp base, which is made of aluminium.
Hirata-san and I will talk about these changes next week.
Other: we installed in ASC0 rack the QPD driver for the POP infrared transmitted beam.
Type-B and Type-Bp TO DO list
- Agree on the changes needed for the Type-Bp cable prototype.
- Provide more information to Miyo-kun for the auto-measurement system.
- For the BS suspension, I calculated the decay time produced by the damper ring of the ~50 mHz yaw mode of the whole suspension: it is 31 s. Therefore, it's not necessary to adjust the position of the damper ring.
- I provided Marc with a Matlab version of the SR2 rigid body model. Back in the day I adjusted the parameters of the model for the calculated transfer functions to match the measured ones. Then I used it to calculate the noise budget that was shown in the Type-B paper.
- Last week, we recieved the cable clamps prototype for the uppermost statge of Type-Bp suspensions. We should test them very soon.
- Hirata-san designed two ways of mounting the limit switches in the traverser. After email discussion with Takahashi-san, Satou-san and me, it seems we'll make prototypes for both version to try.
In PRM, we examined more carefully the condition of H1 channel, whose actuation force seems unusually lower than in other channels. By exchanging cables with other coils, actuating on them sepatrately, and using the oplev TILT_VER channel, we realized that the problem is not within the chamber and it's likely in the coil driver. The effect of H1 on the mirror is at least 50% less of H1 compared with the effect of other coils.
Yokozawa-san reported that, in PRM, the length QPD board broke. Does anyone know the current status?
Type-B and Type-Bp TO DO list
Test Type-Bp cable clamps prototype. This might be a messy task because some cables are entangled inside the chamber and we need to disentangle them. Which suspension should we use?
- Check the cable of the BS F0 fishing rod. There might be something wrong with it.
- I should provide diaggui templates for Type-B automeasurement. Currently, some from Type-A are in place for that.
- If possible, we should visually inspect the BS payload more carefully in case it's close to touching the security structure. After moving it in yaw for green laser alignment, Hirata-san only did a cursory check.
Not for VIS but we were assigned to do it: install the driver for the QPD in POP table.
- BS alignment: Last week on Wednesday, we aligned the BS to reflect the X-arm green laser to reach the Y-End.
We adjusted the yaw of the suspension with the F0 stepper motor. We didn't have much time to check, but Hirata-san did a cursory inspection of the payload and it didn't seem to be touching the security structure. Until now there's no indication there's a problem with the suspension, but we should take a more careful look at it soon. Maybe on a Friday.
During aligmment the oplev went out of range and we recover it on Thursday.
From Wednesday to Thursday the suspension drifted and we had to repeat the X-arm green laser alignment this week again.
As we have talked before, Terrence detected a very large residual motion of the BS optic, so we investigated.
On Thursday last week we removed the temporary duct-shaped wind shields and covered the chamber with our traditional fabric covers. The residual motion dissapear.
Comparison of "before and after" measurements are shown in the klog, including one of the performance of the damper ring, from which I can estimate the decay time of the 50 mHz yaw mode of the chain.
- In PRM, the mirror actuator H1 has a gain which is unusually higher than the others, and we don't know the reason.
There's a klog from September 2019, in which it is suggested the coil driver is not working properly and that it should be replaced.
I superficially assessed the situation by applying actuation in H1, H2 and H4 separately and comparing the motion measured in the oplev QPDs: H1 produced less than half the displacement in yaw than H2 and H4. We definitely should investigate more.
- Hirata-san contined the design of the treverser limit switch holders. He also provided information of the position of the centre of mass of PRM mirror.
Type-B and Type-Bp TO DO list
Replace the PRM mirror coil driver actuator, write the actuation matrix taking into account the position of the CM of the mirror, and estimate the relative imbalance in the actuation channels (e.g. using the Pringle mode). It would be good to calibrate the oplev QPDs and diagonalize the oplev before doig this.
- For the BS magnetic damper, using the data I took last week, calculate the decay time and adjust the position of the ring if necessary.
- Check the cable of the BS F0 fishing rod. There might be something wrong with it.
- We should carry out regular health checks of the suspensions, during night time, for example. We should agree on the implemetation next week.
- BS F1 LVDT:
- After calibration last week, I updated the medm values accordingly, however, the amount of actuation needed to move the keystone to the setpoint was too high (-26,000 counts).
- In order to solve this, I modified the LVDT calibration offset to yield zero, when the keystone is at the nominal 3D-CAD position. I achived this using the old and new calibration data.
- The difference between the setpoints is only 324 um, which is negligible in terms of the uncertainties we have when we measure the height of the optic.
- Ther amount of actuation reduced to around -8,000 counts, and will decrease more when the system goes into vacuum.
- Terrence reported there is some residual motion in the BS suspension. The oplev measures a 50 mHz yaw osciallation which the damper ring is supposed to damp.
- I assessed the performance of the damper ring with the F0 keystone at two different heights, and from visual inspection of the time series, it seems to be working.
- The residual motion is even there with the control system off.
- Currently, we have the impression it might be an enviromental influence. The chamber is not properly shielded against air currents.
- We need to repeat some measuremnts when there is no one working in the tunnel, and if this does not produce good results, we would like to close the chamber with the clean clothes rather than with the improvised aluminium duct.
- Terrence reported the BS oplev was out of range, so we recovered the yaw alignment using the F0 yaw stepper motor. We had to move the length sensing QPD but we took note of the initial and final values in case we want to go back. Terrence measured IM TFs to make sure the payload didn't touch the security structure.
- We assessed the height of the BS:
- We set BF and F0 to zero
- We chose the position of F1 so the optic was at the desired height.
- The amount of actuation F1 required was about -24,000 counts, which is very large, and the keystone moved around -700 um, which is a very small amount given the tools we're using to measure height.
- Given that the reproducibility of the measurement is within 700 um or so (e. g. PRM SF), for the sake of having a low actuation, we could choose a setpoint higher than the one we measured this week.
- We calibrated the BS F1 LVDT using the optical displacement sensor.
- We used a modified holder that allowed us to adjust an additional DoF.
- We realized this device only works with specular reflection, so we had to tilt the sensor a lot (per the manual) and put think ballast mass with a more reflective surface.
- We moved the F1 keystone with coil-magnet actuation only and the range was small. We should use the fishing rod also.
- We got an straight line, which is very good news.
- We should calibrate again with a larger displacement range.
We fixed a traverser cable in PR3. An in-vacuum connector was failing and Hirata-san had to crimp the pins again.
- We finished replacing the flip adapters in SRM
- I compiled the results of the cable health check in one spreadsheet in ordert to compare values that should be similar, and I found two possible problems:
- In SR2, the resistance of IM-V3 coil was higher than expected by about 3 Ohm: we checked and realized the additional resistance comes from a extension cable, but the cable is healthy.
- In the BS, the a resistance of the F0 FR stepper motor is higher by several Ohm, we need to check next week.
- In the BS F1 LVDT, we changed the polrity of the LVDT but we didn't confirmed it worked until this week: positive actuation, which moves the keystone upwards, produces a positive LVDT output.
- We tried to calibrate the BS F1 LVDT today, but we were not able to align the displacement sensor properly. The holder can only be adjusted in one angular DoF, but we need to be able to adjust in teo angular DoF.
I compiled the results of the cable health check and found some potential problems, which we aim to fix next week.
- PR2 BF FR, pins 1-6, P1-7: the resistance is 9.7, which is higher than the values in other BF FRs in PR suspensions (8.8 Ohm maximum).
- PR3 BF V3, pins 1-6, P4-7: the resistance is 13.7 Ohm, which is slightly higher than 11.6 Ohm measured in other LVDTs (approximately).
- PR3 Traverser: there are several resistance values that are several Ohm higher than expected:
- R1, P-2-2: pins 1-2, 1-3, 2-3, 2-6, 3-6, 3-7, 6-7.
- R2, P4-2: pins 1-6, 2-6, 3-6, 6-7.
- Hirata-san and Nakagaki-san removed some unused cables that were hanging close to PR2.
- Hirata-san designed cable clamps to be installed at the upper part Type-Bp suspensions.
- They inspected the PR3 traverser to check whether it had the locking assenmblies or they had been removed.
- We calibrated PRM SF LVDT. The calibration factor hasn't been written in the medm screen yet.
- In PR2, we connected the cable for TM-H4. A cable with the same label has beed connected. There are old cables around that flange and they should be removed.
- Terrence reported SR3 wasn't healthy, the TF measurements looked bad.
- We inspected the suspension and found several security structure screws touching the IM or IRM.
- This suspension was healthy, how did it become unhealthy?
- As they should be, the screws were likley close to the IM and IRM and began touching the payload when yaw was adjusted with the F0 stepper motor.
- When we removed the F0 yaw stepper motor we followed a method to keep the same yaw position, but it seems it was not accurate enough.
- We lost the alignment give by the oplev.
- The incident angle of the green beam might have changed also.
- We fixed a broken connection in IM-H2 OSEM cable. We had to replace the whole connector because the pin was jammed. We used to muiltimeter to check the connections were healthy.
- In PR suspensions there are many connectors that are just hanging. Failure in connections is problem which is likely going to happen again. There are some ideas on how to install cable clamps but we need time and discussion.
- We made a rough test of the optical displacement sensor on the optical table; it seems to be working fine. We plan to calibrate F1 LVDT next week because they won't use the laser.
- In the F1 LVDT adapter box, we exchanged places of the cables connected to pins 2 and 7 in the connector for the cable that goes to the vacuum chamber (primary coil used as a readout coil). This should fix the polarity. We didn't test the modified circuit because of DGS maintenance.
- SR2: we finished changing flip adapters.
- SRM: we changed flip adapters in two flanges.
We glued back the magnets onto the mirror.
- It was not necessary to remove the mid-size baffle. It was possible to do visual inspection of magnets within coil bodies with an inspection mirror.
- Because the RM is not designed to be a positioning jig for the mirror, a concern was the relative alignment of the two bodies.
- The method that we used was to release both bodies and lock them, as much as possible, to the position in which they hung free.
- Despite the absence of two magnets, this method work, likely because the any imbalance affects the RM and mirror in equal amounts becuase they both hang from the IM.
- During the course of this work, we security structie locking screws were withdrawn by defined distances, not like before.
After the gluing, procedure I measured a few transfer functions of the IM and the sysytem seemed to be healthy. However, we should do a more comprenhensive set of measurements. I didn't do it becuase IFO alignment has a higher priority.
During IFO alignment work, we realized the traverser was far away from its centre. Likely, a similar error as PR2 one was made in PRM on Thursday last week.
- At the begining it moved, but later it jammed.
- We realized the problem was that some locking assemblies had not been removed and two of them seemed locked. It seems no other suspension has this locking system.
- We removed them and the traverser now moved smoothly.
- During the course of this work, we security structie locking screws were withdrawn by defined distances, not like before.
It would be good to buy more sligs so we don't have to go back and forth between clean booths to pick them up, please.
- While adjusting the yaw of the suspension with the traverser, the suspension suddently moved a lot in yaw and in pitch also, and the traverser jammed.
- Mechanical problems detected by visual inspection:
- At least two traverser motors reached their mechanical limits.
- Two magnets from the mirror fell off out of the coil body cavity, which was surprising.
- A security structure stud fell off from the ring below the BF onto the suspended breadboard. Clearly, the stud had not been fastened correctly.
- Judging from the position of at least one of the BF damper primary coils, the BF had a lot of tilt.
- On top of the BF there was a cable that was hooking to one of the picomotor assemblies. We don't remember seeing this before, so it migth have been caused by the suspension moving too much in yaw.
- Apparently, a human error caused the problem.
- According to Ikeda-san, the suitable way for us to operate the motors is to override the position of the motor that is stored in the driver.
- He might have forgotten to do this, with the consequence of the driver using an arbitrary value for the current position, thus moving the motor too much.
Another concern would be the velocity, but I didn't find information about this in Ikeda-san's klog. I'll ask him.
- The traverser has some switches, but they are not limit switches, they are for defining the origin of the coordinate system used to measure the position of the stepper motors.
- As described in the klog:
- We locked the suspension.
- Unhook the cable.
- Moved the stepper motors by hand (surprisingly, this was easy to do) until they were centered.
- Released the suspension.